# AI-Enhanced 7-Point Checklist for Melanoma Detection Using Clinical Knowledge Graphs and Data-Driven Quantification

### v1

@@ 5 Discussion
In this study, a novel method to melanoma diagnosis was proposed, integrating clinical empirical knowledge, diagnostic procedures, and graph convolutional networks. Specifically, an appropriate topological graph structure was established by improving the causal learning of attributes associated with melanoma. Computational logic from clinical algorithms was borrowed to enhance explanatory power and clinical acceptance by improving the strong correlation between results and imaging datasets. Moreover, certain elements identified during this study warranted further discussion.
@@
Figure 6: The ROC comparison graph illustrates the performance contrast between the proposed method and the traditional 7PCL algorithm.
Table 5: Comparison of One-Hot and GloVe.
Semantic information Training time AUC
Encoding Total Mel Attributes
One-Hot $\times$ 3s 34 mins 86.8 81.1
GloVe $\checkmark$ 15s 45 mins 89.1 83.7
The results demonstrated the significance of the multi-order node associations introduced by ARFP as topological information, enhancing melanoma detection performance. In preliminary experiments, it was observed that the model’s performance improved as the order $k$ increased from 1 to 3. However, the results plateaued at order $k=4$. Upon closer examination, it was found that as the order $k$ increased, the number of links decreased sharply due to the limited number of nodes (7), resulting in the loss of useful information. Consequently, selecting the optimal order for this work was deemed paramount. Furthermore, the incorporation of additional, more informative labels may potentially enhance the overall outcome.
Additionally, current studies on graph networks related to melanoma detection have primarily focused on link information between nodes. However, for the first time, the semantic information contained in the category names of the nodes themselves was considered. As shown in Table 5, comparison data of predicted knots at the end of data processing time were retained in the study. It was observed that the semantic information of the nodes themselves, while increasing the total time for extraction and processing, resulted in a positive improvement in comparison. Indeed, numerous additional coding methods could be explored in future work.
In presenting the main results of this paper, we used the widely adopted ResNet-50 model to enhance the persuasiveness of our findings. Actually, in the preliminary studies, we experimented with various deep learning architectures, including EfficientNetV2-M, InceptionV3, and MobileNetV2. The results obtained from these models exhibit similar trends and comparisons to those of ResNet-50. The AUC values for melanoma (Mel) across these architectures range from 81.5% to 88.2%. The predictive capacity of our proposed framework may be further enhanced through the utilisation of more sophisticated feature extraction or fusion methodologies, such as the vision transformer (ViT) based models.
6 Conclusion
This study proposes an AI-enhanced 7PCL for melanoma detection using a unique topology enriched with clinical insights, simulating dermatologists’ diagnostic behavior. By considering directed multi-order relationships and a gradient prediction structure, our method provides weight measures for each attribute. This helps dermatologists intuitively analyze attributes, enhancing interpretability in clinical scenarios and improving detection performance and acceptance.
[…]

### v3

@@ VI Discussion
In this study, we proposed a novel approach to melanoma diagnosis, integrating clinical empirical knowledge, diagnostic procedures, and GCNs. Specifically, we improved the causal learning of melanoma-associated attributes to construct an optimized topological graph structure. By incorporating computational logic from clinical algorithms, our method enhances interpretability and clinical acceptance by strengthening the correlation between diagnostic outcomes and imaging datasets. Additionally, several key aspects identified during this study warrant further discussion.
@@
Figure 8: ROC comparison graph showcasing the performance differences between the proposed method and the traditional 7PCL algorithm across two classification tasks.(a): Classification of Mel and Nevi; (b): Classification of Mel and other diseases (MISC, Nevi, BCC, and SK)
VI-A Multi-Order Node Associations and Their Impact
Our results demonstrate that multi-order node associations, introduced through ARFP, play a crucial role in melanoma detection performance. Beyond the main results presented in the results section, additional experimental evaluations reveal a consistent improvement in AUC as the order $k$ increases: 87.1% at $k=1$, 90.6% at $k=2$, and 92.0% at $k=3$, indicating that expanding the receptive field effectively captures more informative attribute relationships. However, at $k=4$, AUC slightly declines to 91.8%, suggesting a performance plateau. Given that the 7PCL graph comprises only 7 nodes, further increasing $k$ reduces the number of meaningful paths, leading to information dilution rather than enrichment.
These observations highlight the importance of selecting an optimal order $k$ to balance local feature aggregation and global information propagation. Future work may explore dynamic receptive field adjustment or adaptive graph construction based on dataset-specific connectivity patterns. Additionally, incorporating more informative labels could help mitigate the limitations imposed by the small number of nodes, potentially enhancing graph-based melanoma classification performance.
VI-B Semantic Information and Its Role in Graph Learning
Existing research on graph-based melanoma detection has primarily focused on link relationships between nodes, while the semantic content of node labels has often been overlooked [43, 22, 38]. To the best of our knowledge, this study is the first to integrate dynamic semantic information into both feature extraction and the graph learning process, enabling full utilization of diagnostic knowledge.
As shown in Table V, we evaluated multiple semantic embedding methods, including one-hot encoding [46], GloVe [47], Word2Vec [48], and BERT [37]. Each method differs in its ability to capture contextual relationships among node labels, with dynamic embeddings offering greater adaptability.
Results indicate that models incorporating semantic embeddings outperform those without them, with BERT achieving the highest AUC values of 92.0% for melanoma classification and an average AUC of 88.6% across all attributes. However, the inclusion of dynamic embeddings comes at the cost of increased computational time, particularly with BERT requiring 84 minutes of training time compared to 56-63 minutes for static embeddings. Despite this trade-off, the improved diagnostic performance demonstrates the effectiveness of integrating contextual knowledge into graph-based melanoma detection.
These findings suggest that leveraging dynamic embeddings allows the model to capture higher-order semantic relationships, ultimately enhancing classification performance. Future research could explore hybrid approaches, balancing computational efficiency with embedding effectiveness for large-scale dermatological datasets.
TABLE V: Comparison of Different Semantic Embedding Methods.
Semantic Information Dynamic Embedding Training time AUC
Encoding Total Mel Avg
One-Hot $\times$ $\times$ 0.3s 63 mins 86.8 81.1
GloVe $\checkmark$ $\times$ 1s 61 mins 89.1 83.7
Word2Vec $\checkmark$ $\times$ 1s 56 mins 88.7 84.0
BERT $\checkmark$ $\checkmark$ 3s 84 mins 92.0 88.6
VI-C Clinical Implications and Adaptive Weighting for 7PCL
From a clinical perspective, the 7PCL algorithm was originally developed to assist dermatologists in distinguishing Mel from Nevi based on dermoscopic features. However, adapting it into a modern computer-aided diagnosis (CAD) system that encompasses a broader range of similar-looking skin lesions and integrates both clinical and dermoscopic features presents significant challenges. One key limitation is that the original scoring system, which assigns a weight of 2 to major features and 1 to minor features, was designed specifically for Mel and Nevi differentiation and may not generalize well to more diverse diagnostic scenarios [7, 49, 50].
To address these limitations, this study introduces a data-driven approach that dynamically adjusts attribute weights based on the classification task. By refining the weighting mechanism, the 7PCL algorithm can adapt to different dermatological conditions while maintaining its clinical interpretability. This adjustment allows the system to better align with real-world diagnostic needs, making it more applicable in various clinical settings and improving its utility as a decision-support tool for medical professionals.
VI-D Potential for Integrating Foundation Models
Foundation models, particularly large-scale multimodal architectures, have demonstrated remarkable adaptability across various domains, including dermatology. These models, pretrained on vast and diverse datasets, offer strong feature representations that can be fine-tuned for specific medical tasks, including melanoma diagnosis. Given their ability to capture complex hierarchical patterns and cross-modal interactions, integrating foundation models into our framework could further enhance its robustness and generalization. Specifically, vision-language models like SkinGPT-4 and pathology foundation models have shown promise in leveraging both textual clinical knowledge and visual cues, potentially enriching the diagnostic reasoning process in automated systems [51, 52, 53, 54].
Incorporating foundation models into our approach could provide several advantages. First, they could refine the dynamic weighting mechanism by learning from broader datasets, improving the adaptability of the 7PCL-based scoring system beyond its original scope. Second, their self-supervised learning capabilities could mitigate the limitations of annotated medical datasets, allowing for improved feature extraction even in data-constrained environments. Lastly, by integrating pretrained multimodal embeddings, our framework could benefit from enhanced attribute interactions and context-aware diagnostics, making it more interpretable and clinically applicable. Future work will explore the feasibility of integrating foundation models to further advance automated melanoma diagnosis.
VII Conclusion
In this study, we present an artificial intelligence-enhanced 7-Point Checklist (7PCL) method for melanoma detection, designed to integrate a clinically-informed topology that mirrors the diagnostic approach of dermatologists. This approach addresses the limitations of the traditional 7PCL, which has been primarily restricted to differentiating Mel from Nevi, by enabling automated adaptation to diverse diagnostic scenarios. Our method efficiently extracts visual features using an architecture that combines single-modal and cross-modal attention mechanisms, capturing subtle relationships within and across modalities. By incorporating directed multi-order relationships and gradient prediction structures, our approach dynamically assigns weighted metrics to each attribute. This allows the algorithm to adapt to varying diagnostic tasks, such as distinguishing Mel from a broader range of skin conditions with overlapping visual characteristics. This enhanced weighting system not only improves detection performance but also provides dermatologists with a more intuitive and interpretable framework for attribute analysis. By aligning the algorithm’s design with clinical workflows, the proposed method could enhance both the effectiveness and acceptance of AI-assisted diagnosis in real-world settings.